OpenStudy (pujan20):

hi i need help for alg 1 honors

OpenStudy (anonymous):

ok whats the question

OpenStudy (pujan20):

3x+2 _______ 8

OpenStudy (pujan20):

i need to find the sum equivilent

OpenStudy (laughoutloud):

the sum equivalent to what?

OpenStudy (pujan20):

the expression.

OpenStudy (laughoutloud):

what are your answer choices?

OpenStudy (laughoutloud):

if any?

OpenStudy (pujan20):

no choices. like my alg. 1 teacher is seeing our progress, but its 20% smh

OpenStudy (laughoutloud):

does the expression look like this? \[\frac{ 3x+2 }{ 8 }\]

OpenStudy (pujan20):

yep sorry, I'm used to typing equations on mathway

OpenStudy (pujan20):

doesn't say. the question is what sum or difference is equivalent to the expression

OpenStudy (laughoutloud):

that doesn't make sense?

OpenStudy (pujan20):

ik. I'm usually good at algebra

OpenStudy (laughoutloud):

same it's my favorite subject

OpenStudy (anonymous):

I think this might be asking you to split this into 2 fractions

OpenStudy (laughoutloud):

I was thinking the same but the wording is off

OpenStudy (pujan20):

i think its

OpenStudy (theloshua):

no, haham this is easy. theyre asking you to solve for x

OpenStudy (pujan20):

\[ \frac{ 5x }{ 8 }\]

OpenStudy (laughoutloud):

you can't solve there is no equal sign

OpenStudy (theloshua):

no, you cannot add the 2 to the 3x because they are different variables

OpenStudy (pujan20):

ohh

OpenStudy (anonymous):

no. 3x and 2 aren't like terms, so they can't be combined \[\frac{ 3x+2 }{ 8 }=\frac{ 3x }{ 8 }+\frac{ 2 }{ 8 }\] then reduce 2/8

OpenStudy (theloshua):

and yes, you can solve for it :)

OpenStudy (laughoutloud):

how?

OpenStudy (pujan20):

but I'm supposed to find the equivalent sum

OpenStudy (pujan20):

instead of solving it itself

OpenStudy (theloshua):

hold on, let me think. Im trying to figure it out. im pretty sure this has to do with multiplying the fraction

OpenStudy (anonymous):

3x/8 + 2/8 is an equivalent sum to (3x + 2)/8

OpenStudy (pujan20):

omg how didn't i realize that!

OpenStudy (laughoutloud):

yes peachpi is right now you just have to reduce the 2/8

OpenStudy (pujan20):

into 1/4

OpenStudy (laughoutloud):

yes

OpenStudy (pujan20):

thx, but i have 1 more problem like this

OpenStudy (pujan20):

can u help

OpenStudy (laughoutloud):

I can try?

OpenStudy (pujan20):

\[\frac{ 3x-2 }{ 9}\]

OpenStudy (laughoutloud):

k so split it into two again

OpenStudy (pujan20):

so 3x/9 - 2/9?

OpenStudy (laughoutloud):

yes but the 3x/9 can reduce... I think @peachpi

OpenStudy (anonymous):

yes it does reduce

OpenStudy (pujan20):

into 1x/3 maybe

OpenStudy (laughoutloud):

yes

OpenStudy (pujan20):

omg thx so much!

OpenStudy (pujan20):

u all deserve best response ^v^

OpenStudy (laughoutloud):

your welcome